OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? -b b^2 - 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0 Answer

OpenStudy (anonymous):

b² - 4ac

OpenStudy (anonymous):

(7)^2 -4(2)(3)=25 which is perfect square therefore you can use quadratic equation

OpenStudy (anonymous):

@Brad1996 , xD

OpenStudy (anonymous):

2x^2 + 7x + 3 = 0 (2x+1)(x+3)=0 x1=-1/2,x2=-3

OpenStudy (anonymous):

wait whats the answer I got 2 diff answers :l

OpenStudy (anonymous):

mine , because you said to Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? -b b^2 - 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0

OpenStudy (anonymous):

lily just factor it

OpenStudy (anonymous):

@Brad1996 right xD

OpenStudy (anonymous):

so its 2x^2 + 7x + 3 = 0?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

2x^2 + 7x + 3 = 0 b^2 - 4ac where b=7, a=2, c=3 (7)^2 -4(2)(3)=25

OpenStudy (anonymous):

so the answer is 7^2 -4(2)(3)=25?

OpenStudy (anonymous):

yup 25 , \ man xD